Solve the quadratic equation 12x^2 - 3x = 0 by factoring
Answer by stanbon(75887)

Factor to get (3x)(4x-1)=0
Since the product is zero one of the
factors must be zero.
3x=0 or 4x-1=0
x=0 or x=1/4
Cheers,
Stan H.